You are given the following cost data:Total fixed costs are $60.


(q,TVC)➡️ (1,25) (2,40) (3,60) (4,90) (5,130) (6,185) .How many units of output will this firm produce at a price of $22 and at a price of $42? What is the total revenue and total cost at each price? What is the profit at each price? Briefly explain using the concept of marginal cost.

1
Expert's answer
2022-01-19T05:59:21-0500



P>MC, P>AVC, P>ATC

MR=MC Profit maximization


1)P=22

Q=3

TR=66

TC=120

TR-TC=66-120=-54

2) P=42

Q=6

TR=252

TC=245

TR-TC=252-245=7


In the first case a profit, in the second case a loss